OpenStudy (anonymous):

A sequence is defined recursively using the equation f(n + 1) = f(n) – 8. If f(1) = 100, what is f(6)?

OpenStudy (loser66):

find f2, f3, f4 to get f5

OpenStudy (kainui):

This is basically a differential equation with initial conditions... Err, telescoping series, \[\sum_{n=1}^5 f(n+1)-f(n)=\sum_{n=1}^5 -8 \] \[f(6)-f(1)=-8*5\]